35 35 votes For a given biased coin, the probability that the outcome of a toss is a head is $0.4$. This coin is tossed $1,000$ times. Let $X$ denote the random variable whose value is the number of times that head appeared in these $1,000$ tosses. The standard deviation of $X$ (rounded to $2$ decimal place) is _________ Probability gatecse-2021-set2 numerical-answers probability random-variable one-mark statistics + – Arjun 13.7k views answer comment Share Follow Print See 1 comment 1 1 comment reply KhushalX commented Feb 8 reply Follow flag Here, X is a binomial random variable. For X, variance is n*p*(1-p). Here, p is probability of getting a heads i.e. p = 0.4. So, variance = 1000 * 0.4 * 0.6 Hence, standard deviation = square root of variance = 15.49 (approx to two decimal places) 0 0 replyShare Please log in or register to add a comment.
Best answer 47 47 votes It is Independent trails with same rate of success. So we can use Binomial distribution Probability of success $(p) = 0.4$ Probability of failure $(q) = 1 – 0.4 = 0.6$ Number of trials $(n) = 1000$ Variance $=n\times p\times q = 1000 * 0.4 * 0.6 = 240$ Standard deviation $=\sqrt{240} = 15.49$ zxy123 answered Feb 18, 2021 • edited Apr 4, 2021 by gatecse zxy123 comment Share Follow See all 5 Comments 5 5 Comments reply Show 2 previous comments Santi Singha commented Dec 20, 2024 reply Follow flag Why we don't use Poisson distribution? λ = np = 1000 * 0.4 = 400 In Poisson, mean = variance = λ So variance = 400 and standard deviation = √400 = 20 0 0 replyShare Kshitij Sharma commented Dec 25, 2024 reply Follow flag @Santi SinghaNormally, Poisson distribution ka use more suitable hota hai jab p kafi chhota hota hai (jaise 0.01 ya usse kam).Binomial distribution ka use tab hota hai jab:Fixed number of trial hote hai.Har trial me sirf do outcomes possible hote hain (success ya failure).Har trial independent hota hai.Success ki probability p har trial me same hota hai.Poisson DistributionPoisson distribution typically use hota hai jab:Keywords dikhe : jaise accidents per day, calls per hour, etc.Ye "rare events" ke liye use kiya jata hai jab n bahut bada hota hai aur p bahut chhota hota hai. 28 28 replyShare firstgod commented Mar 27 reply Follow flag we cant because in poissons the probab. of success is very very small 0 0 replyShare Please log in or register to add a comment.
0 0 votes Good Luck soudipta_dutta answered Feb 7 soudipta_dutta comment Share Follow 0 reply Please log in or register to add a comment.